# CMSO-transducing tree-like graph decompositions

### v2

@@ 2.2 Logic and transductions
Given a $\Sigma$-structure $\mathbb{A}$ and, for some vocabulary $\Gamma$, a $\Gamma$-structure $\mathbb{B}$, we write $\mathbb{A}\sqsubseteq\mathbb{B}$ if $\Sigma\subseteq\Gamma$, $U_{\mathbb{A}}\subseteq U_{\mathbb{B}}$ and for each symbol $Q$ in $\Sigma$, $Q_{\mathbb{A}}=Q_{\mathbb{B}}$. [^3] We require equality here (in particular only elements or subsets of $U_{\mathbb{A}}$ are related in $Q_{\mathbb{B}}$). This differs from classical notions of inclusions of relational structures which typically require equality only on the restriction of the universe to $U_{\mathbb{A}}$, i.e., $Q_{\mathbb{A}}=Q_{\mathbb{B}_{/U_{\mathbb{A}}}}$, e.g., in order to correspond to induced graphs. We write $\mathbb{A}\sqcup\mathbb{B}$ to denote the $(\Sigma\cup\Gamma)$-structure consisting of the universe $U_{\mathbb{A}}\cup U_{\mathbb{B}}$ and, for each symbol $Q\in\Sigma\cup\Gamma$, the interpretation $Q_{\mathbb{A}\sqcup\mathbb{B}}$ which is $Q_{\mathbb{A}}$, $Q_{\mathbb{B}}$, or $Q_{\mathbb{A}}\cup Q_{\mathbb{B}}$ according to whether $Q$ belongs to $\Sigma\setminus\Gamma$, to $\Gamma\setminus\Sigma$, or to $\Sigma\cap\Gamma$ [^4] We do not require $\mathbb{A}$ and $\mathbb{B}$ to be disjoint structures whence we may have $\mathbb{A}\not\sqsubseteq\mathbb{A}\sqcup\mathbb{B}$.
To describe properties of (extended) relational structures, we use monadic second order logic ($\operatorname{MSO}$) and refer for instance to [15, 20, 23, 27] for the definition of $\operatorname{MSO}$ on extended relational structures such as matroids or set systems in general. This logic allows quantification both over single elements of the universe and over subsets of the universe. We also use counting $\operatorname{MSO}$ ($\operatorname{C}{\!\!}\operatorname{MSO}$), which is the extension of $\operatorname{MSO}$ with, for every positive integer $p$, a unary set predicate $\operatorname{C}_{p}$ that checks whether the size of a subset is divisible by $p$ or not. We only use $\operatorname{C}_{2}$. As usual, lowercase variables indicates first-order-quantified variables, while uppercase variables indicates monadic-quantified variables. For a formula $\phi$, we write, e.g.$\phi(x,y,X)$ to indicate that the variables $x$, $y$, and $X$ belong to the set of free variables of $\phi$, namely, the set of variables occurring in $\phi$ that are not bound to a quantifier within $\phi$. A sentence is a formula without free variables.
@@
Let $\Sigma$ and $\Gamma$ be two extended vocabularies. A $\Sigma$-to-$\Gamma$ transduction is a set $\tau$ of pairs formed by a $\Sigma$-structure, call the input, and a $\Gamma$-structure, called the output. We write $\mathbb{B}\in\tau(\mathbb{A})$ when $(\mathbb{A},\mathbb{B})\in\tau$. When for every pair $(\mathbb{A},\mathbb{B})\in\tau$ we have $\mathbb{A}\sqsubseteq\mathbb{B}$, we call $\tau$ an overlay transduction. Some transductions can be defined by means of $\operatorname{MSO}$- or $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formulas. This leads to the notion of $\operatorname{MSO}$- and $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions. Following the presentation of [4], for $\mathsf{L}$ denoting $\operatorname{MSO}$ or $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$, define an $\mathsf{L}$-transduction to be a transduction obtained by composing a finite number of atomic $\mathsf{L}$-transductions of the following kinds.
Filtering.
@@
Universe restriction.
A transduction specified by a $\mathsf{L}$-formula $\phi$ over the input vocabulary $\Sigma$, with one free first-order variable, which restricts the universe to those elements that satisfy $\phi$. The output vocabulary is $\Sigma$ and the interpretation of every relation (resp. every predicate) in the output structure is defined as the restriction of its interpretation in the input structure to those tuples of elements satisfying $\phi$ (resp. tuples of sets of elements that satisfy $\phi$). This defines a total function from $\Sigma$-structures to $\Sigma$-structures.
Interpretation.
@@
Copying.
An overlay transduction parametrized by a positive integer $k$ that adds $k$ copies of each element to the universe. The output vocabulary consists in the input vocabulary $\Sigma$ extended with $k$ binary relational symbols $(\mathsf{copy}_{i})_{i\in[k]}$ interpreted as pairs of elements $(x,y)$ saying that “$y$ is the $i$-th copy of $x$”. The interpretation of the relations (resp. predicates) of the input structure are preserved, on original elements. This defines a total function from $\Sigma$-structures to $\Gamma$-structures, where $\Gamma=\Sigma\cup\{\mathsf{copy}_{i}\mid i\in[k]\}$.
Colouring.
An overlay transduction that adds a new unary relation $\mathsf{color}\notin\Sigma$ to the structure. Any possible interpretation yields an output; indeed the interpretation is chosen non-deterministically. The interpretation of the relations (resp. predicates) of the input structure are preserved. Hence, it defines a total (non-functional) relation from $\Sigma$-structures to $\Gamma$-structures where $\Gamma=\Sigma\cup\{\mathsf{color}\}$ (providing $\mathsf{color}\notin\Sigma$).
We say an $\mathsf{L}$-transduction is deterministic if it does not use colouring, it is non-deterministic otherwise. By definition, deterministic $\mathsf{L}$-transductions define functions.
3 Transducing the laminar-tree
@@
Theorem 3.1.
Let $\Sigma$ be an extended vocabulary, including a unary set predicate name $\mathsf{SET}$ and not including the binary relational symbol $\mathsf{ancestor}$. There exists a non-deterministic $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each laminar set system $(U,\mathcal{S})$ represented as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ and inducing the laminar tree $T$ with ${L(T)=U}$, and for each $\Sigma$-structure $\mathbb{A}$ with $\mathbb{S}\sqsubseteq\mathbb{A}$, $\tau(\mathbb{A})$ is non-empty and every output in $\tau(\mathbb{A})$ is equal to some $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ representing $T$.
@@
We prove Theorem 3.1 in Section 3.2, using the key tools developed in Section 3.1, that allow us to represent each inner node of $T$ with a pair of leaves from its subtree, while keeping the number of inner nodes represented by a given leaf bounded.
3.1 Inner node representatives
@@
Let $S\subseteq V\setminus L$, and let $(\pi,\sigma)$ be a pair of injective mappings from $S$ to $L$. We say that the pair $(\pi,\sigma)$ identifies $S$ if for each $s\in S$, $s$ is the least common ancestor of $\pi(s)$ and $\sigma(s)$. For $s\in S$ and $x\in V$, we say that $x$ is $s$-requested in $(\pi,\sigma)$ if $x$ lies on the path from $\pi(s)$ to $\sigma(s)$ (namely, on either of the paths from $\pi(s)$ to $s$ and from $\sigma(s)$ to $s$). We say that $x$ is requested in $(\pi,\sigma)$ if it is $s$-requested for some $s$. The pair $(\pi,\sigma)$ has unique request if every node $x$ of $T$ is requested at most once in $(\pi,\sigma)$, i.e., there exists at most one $s\in S$ such that $x$ is $s$-requested in $(\pi,\sigma)$. Note that if $(\pi,\sigma)$ has unique request then the paths from $\pi(s)$ to $\sigma(s)$ are pairwise disjoint for all the $s\in S$. We now state a few basic observations.
Remark 3.2.
@@
  3. 3.
    For each $s\in S$, $s$ is $s$-requested in $(\pi,\sigma)$.
  4. 4.
@@
Lemma 3.3.
Let $(\pi,\sigma)$ identifying some subset $S$ of $V\setminus L$ with unique request. Then for each $a\in\pi(S)$ the node $\pi^{-1}(a)$ is the least ancestor of $a$ which belongs to $S$.
Proof 3.4.
Let $a\in\pi(S)$ and let $s=\pi^{-1}(a)$. By definition, $s$ is an ancestor of $a$ which belongs to $S$. Let $y$ be the least ancestor of $a$ that is contained in $S$. As $s$ is an ancestor of $a$ belonging to $S$, $y$ must be a descendant of $s$. Hence, $y$ is $s$-requested in $(\pi,\sigma)$. Additionally, by Item 3 of Remark 3.2, $y$ is $y$-requested in $(\pi,\sigma)$. Since $(\pi,\sigma)$ has unique request, $y=s$. Thus, $s$ is the least ancestor of $a$ which belongs to $S$.
Notice that, by Item 1 of Remark 3.2, a similar result holds for each $b\in\sigma(S)$. It follows that the sets $\pi(S)$ and $\sigma(S)$ characterize $(\pi,\sigma)$.
Lemma 3.5.
Let $S\subseteq V\setminus L$, and $(\pi,\sigma)$ and $(\pi^{\prime},\sigma^{\prime})$ be two pairs of injections from $S$ to $L$ identifying $S$ with unique request. If $\pi(S)=\pi^{\prime}(S)$ and $\sigma(S)=\sigma^{\prime}(S)$ then $\pi=\pi^{\prime}$ and $\sigma=\sigma^{\prime}$.
Proof 3.6.
Let $s\in S$, $a=\pi(s)$, and $s^{\prime}=\pi^{\prime-1}(a)$. Both $s$ and $s^{\prime}$ are the least ancestor of $a$ which belongs to $S$, hence $s=s^{\prime}$ and $\pi^{\prime}(s)=a$. Thus $\pi=\pi^{\prime}$. By Item 1 of Remark 3.2, we also obtain that $\sigma=\sigma^{\prime}$.
Let $A$ and $B$ be two disjoint subsets of $L$. We call such a pair $(A,B)$ a bi-colouring. We say that $(A,B)$ identifies $S$ if there exists a pair $(\pi,\sigma)$ identifying $S$ with unique request such that $\pi(S)=A$ and $\sigma(S)=B$. By the previous lemma, for a fixed set $S\subseteq V\setminus L$ the pair $(\pi,\sigma)$ identifying $S$ is unique when it exists. We will also prove that $S$ is actually uniquely determined from $(A,B)$. Before that, we state the following technical lemma.
Lemma 3.7.
@@
  1. 1.
    $x\notin S$, $x$ is not requested in $(\pi,\sigma)$, and for each child $c$ of $x$, $\left|A\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$;
  2. 2.
    $x\notin S$, $x$ is requested in $(\pi,\sigma)$, and there exists one leaf $z\in(A\cup B)\cap V(T_{x})$ such that, for each child $c$ of $x$, $\left|(A\setminus\{z\})\cap V(T_{c})\right|=\left|(B\setminus\{z\})\cap V(T_{c})\right|$;
  3. 3.
    $x\in S$, $x$ is requested in $(\pi,\sigma)$, and there exists two leaves $a\in A\cap V(T_{x})$ and $b\in B\cap V(T_{x})$ such that, for each child $c$ of $x$, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|(B\setminus\{b\}\cap V(T_{c}))\right|$ and $\{a,b\}\nsubseteq V(T_{c})$.
Proof 3.8.
Let $x$ be an inner node. We consider the set $S^{\prime}=S\setminus(V(T_{x})\setminus\{x\})$ of all nodes which are not proper descendants of $x$ and the restrictions $\pi^{\prime}$ and $\sigma^{\prime}$ of, respectively, $\pi$ and $\sigma$ to $S^{\prime}$. By Item 2 of Remark 3.2 $(\pi^{\prime},\sigma^{\prime})$ identify $S^{\prime}$ with unique request. We denote $A^{\prime}=\pi^{\prime}(S^{\prime})$ and $B^{\prime}=\sigma^{\prime}(S^{\prime})$, thus $(A^{\prime},B^{\prime})$ identify $S^{\prime}$. Let $A^{\prime}_{x}=A^{\prime}\cap V(T_{x})$ and $B^{\prime}_{x}=B^{\prime}\cap V(T_{x})$ be the sets of representatives contained in $T_{x}$ of nodes in $S^{\prime}$. Let $a$ be an element of $A^{\prime}_{x}$. The element $s_{a}=\pi^{-1}(a)$ is an ancestor of $x$ because it is an ancestor of $a\in V(T_{x})$ and belongs to $S^{\prime}$ whence not to $V(T_{x})\setminus\{x\}$. Therefore, $x$ is $s_{a}$-requested. As $(\pi^{\prime},\sigma^{\prime})$ has unique request, there exists at most one element in $A^{\prime}_{x}$. Similarly, $B^{\prime}_{x}$ has size at most $1$. Moreover, $A^{\prime}_{x}\cap B^{\prime}_{x}=\emptyset$ by Item 4 of Remark 3.2. We thus we have three cases:
Case $A^{\prime}_{x}\cup B^{\prime}_{x}=\emptyset$.
@@
Case $A^{\prime}_{x}\cup B^{\prime}_{x}=\{a\}$.
@@
Let $c$ be a child of $x$. If $\left|A\cap V(T_{c})\right|\neq\left|B\cap V(T_{c})\right|$, then it means that there exists $z\in(A\cup B)\cap V(T_{c})$, such that either $z\in A$ and $s_{z}=\pi^{-1}(z)\notin V(T_{c})$, or $z\in B$ and $s_{z}=\sigma^{-1}(z)\notin V(T_{c})$. In both cases, $s_{z}$ is a proper ancestor of $c$, whence an ancestor of $x$. Thus, $x$ is $s_{z}$-requested in $(\pi,\sigma)$. However, $x$ is $s_{a}$-requested in $(\pi,\sigma)$ which has unique request, hence $s_{z}=s_{a}$. If $z\in B$, it follows that $z=b$, which contradicts the fact $b\notin V(T_{x})$. Hence, $z\in A$ and thus, $z=\pi(s_{a})=a$. Therefore, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$ for each child $c$ of $x$.
Case $A^{\prime}_{x}=\{a\}$ and $B^{\prime}_{x}=\{b\}$.
Let $s_{a}=\pi^{-1}(a)$ and $s_{b}=\sigma^{-1}(b)$. The node $x$ is $s_{a}$-requested and $s_{b}$-requested in $(\pi,\sigma)$, so, by the unique request property, $s_{a}=s_{b}$. Since $s_{a}$ is the least common ancestor of $a$ and $b$, it belongs to $V(T_{x})$ whence $s_{a}=x$ implying $x\in X$.
Let $c$ be a child of $x$. If $\left|A\cap V(T_{c})\right|\neq\left|B\cap V(T_{c})\right|$, then there exists $z\in(A\cup B)\cap V(T_{c})$ such that either $z\in A$ and $s_{z}=\pi^{-1}(z)\notin V(T_{c})$, or $z\in B$ and $s_{z}=\sigma^{-1}(z)\notin V(T_{c})$. In both cases, $s_{z}$ is a proper ancestor of $c$, whence an ancestor of $x$, and thus $x$ is $s_{z}$-requested in $(\pi,\sigma)$. However, $x$ is $x$-requested in $(\pi,\sigma)$ which has unique request, hence $s_{z}=x$ and thus $z\in\{a,b\}$. Therefore, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|(B\setminus\{b\})\cap V(T_{c})\right|$ for each child $c$ of $x$. Because the least common ancestor of $a$ and $b$ is $x$, there is no child $c$ of $x$ containing both $a$ and $b$ as leaves, i.e., $\{a,b\}\nsubseteq V(T_{c})$.

### v4

@@ 2.2. Logic and transductions
Given a $\Sigma$-structure $\mathbb{A}$ and, for some vocabulary $\Gamma$, a $\Gamma$-structure $\mathbb{B}$, we write $\mathbb{A}\sqsubseteq\mathbb{B}$ if $\Sigma\subseteq\Gamma$, $U_{\mathbb{A}}\subseteq U_{\mathbb{B}}$ and for each symbol $Q$ in $\Sigma$, $Q_{\mathbb{A}}=Q_{\mathbb{B}}$. [^3] We require equality here (in particular only elements or subsets of $U_{\mathbb{A}}$ are related in $Q_{\mathbb{B}}$). This differs from classical notions of inclusions of relational structures which typically require equality only on the restriction of the universe to $U_{\mathbb{A}}$, i.e., $Q_{\mathbb{A}}=Q_{\mathbb{B}_{/U_{\mathbb{A}}}}$, e.g., in order to correspond to induced graphs. We write $\mathbb{A}\sqcup\mathbb{B}$ to denote the $(\Sigma\cup\Gamma)$-structure consisting of the universe $U_{\mathbb{A}}\cup U_{\mathbb{B}}$ and, for each symbol $Q\in\Sigma\cup\Gamma$, the interpretation $Q_{\mathbb{A}\sqcup\mathbb{B}}$ which is $Q_{\mathbb{A}}$, $Q_{\mathbb{B}}$, or $Q_{\mathbb{A}}\cup Q_{\mathbb{B}}$ according to whether $Q$ belongs to $\Sigma\setminus\Gamma$, to $\Gamma\setminus\Sigma$, or to $\Sigma\cap\Gamma$. [^4] We do not require $\mathbb{A}$ and $\mathbb{B}$ to be disjoint structures whence we may have $\mathbb{A}\not\sqsubseteq\mathbb{A}\sqcup\mathbb{B}$.
[…]
@@
Let $\Sigma$ and $\Gamma$ be two extended vocabularies. A $\Sigma$-to-$\Gamma$ transduction is a set $\tau$ of pairs formed by a $\Sigma$-structure, call the input, and a $\Gamma$-structure, called the output. We write $\mathbb{B}\in\tau(\mathbb{A})$ when $(\mathbb{A},\mathbb{B})\in\tau$. We say that a class of $\Sigma$-structures $\mathcal{C}$ transduces a class of $\Gamma$-structures $\mathcal{D}$ if there is a $\Sigma$-to-$\Gamma$ transduction $\tau$ with $\mathcal{D}\subseteq\tau(\mathcal{C})$. When for every pair $(\mathbb{A},\mathbb{B})\in\tau$ we have $\mathbb{A}\sqsubseteq\mathbb{B}$, we call $\tau$ an overlay transduction. Overlay transductions are used to augment a relational structure by adding additional relations while not altering existing relations in any way. Some transductions can be defined by means of $\operatorname{MSO}$- or $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-formulas. This leads to the notion of $\operatorname{MSO}$- and $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transductions. Following the presentation of [BGP21], for $\mathsf{L}$ denoting $\operatorname{MSO}$ or $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$, define an $\mathsf{L}$-transduction to be a transduction obtained by composing a finite number of atomic $\mathsf{L}$-transductions of the following kinds.
Filtering:
@@
Universe restriction:
A transduction specified by an $\mathsf{L}$-formula $\phi$ over the input vocabulary $\Sigma$, with one free first-order variable, which restricts the universe to those elements that satisfy $\phi$. The output vocabulary is $\Sigma$ and the interpretation of every relation (resp. every predicate) in the output structure is defined as the restriction of its interpretation in the input structure to those tuples of elements satisfying $\phi$ (resp. tuples of sets of elements that satisfy $\phi$). This defines a total function from $\Sigma$-structures to $\Sigma$-structures.
Interpretation:
@@
Copying:
An overlay transduction parametrized by a positive integer $k$ that adds $k$ copies of each element to the universe. The output vocabulary consists in the input vocabulary $\Sigma$ extended with $k$ binary relational symbols $(\mathsf{copy}_{i})_{i\in[k]}$ interpreted as pairs of elements $(x,y)$ saying that “$y$ is the $i$-th copy of xx”. The interpretation of the relations (resp. predicates) of the input structure are preserved, on original elements. This defines a total function from $\Sigma$-structures to $\Gamma$-structures, where $\Gamma=\Sigma\cup\{\mathsf{copy}_{i}\mid i\in[k]\}$.
Colouring:
An overlay transduction that adds a new unary relation $\mathsf{color}\notin\Sigma$ to the structure. Any possible interpretation yields an output; indeed the interpretation is chosen non-deterministically. The interpretation of the relations (resp. predicates) of the input structure are preserved. Hence, it defines a total (non-functional) relation from $\Sigma$-structures to $\Gamma$-structures where $\Gamma=\Sigma\cup\{\mathsf{color}\}$. Note that by composing several transductions of this type any given element may receive multiple colors.
We say an $\mathsf{L}$-transduction is deterministic if it does not use colouring, it is non-deterministic otherwise. By definition, deterministic $\mathsf{L}$-transductions define partial functions. It is known that every $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction can be put in a standard form in which each of the above atomic transductions occurs once in a fixed order, namely: first colouring, then filtering, copying, interpreting the output relations, and finally restricting the universe [CE12].
3. Transducing the laminar-tree
@@
Theorem 3.
Let $\Sigma$ be an extended vocabulary, including a unary set predicate name $\mathsf{SET}$ and not including the binary relational symbol $\mathsf{ancestor}$. There exists a non-deterministic overlay $\operatorname{C}_{2}{\!\!}\operatorname{MSO}$-transduction $\tau$ such that, for each laminar set system $(U,\mathcal{S})$ represented as the $\{\mathsf{SET}\}$-structure $\mathbb{S}$ and inducing the laminar tree $T$ with L(T)=U{L(T)=U}, and for each $\Sigma$-structure $\mathbb{A}$ with $\mathbb{S}\sqsubseteq\mathbb{A}$, $\tau(\mathbb{A})$ is non-empty and every output in $\tau(\mathbb{A})$ is equal to $\mathbb{A}\sqcup\mathbb{T}$ for some $\{\mathsf{ancestor}\}$-structure $\mathbb{T}$ representing $T$.
@@
We prove 3 in Section 3.2, using the key tools developed in Section 3.1, that allow us to represent each inner node of $T$ with a pair of leaves from its subtree, while keeping, for each leaf, the number of inner nodes it represents bounded.
3.1. Inner node representatives
@@
Let $S\subseteq V\setminus L$ be a set of inner nodes, and let $(\pi,\sigma)$ be a pair of injective mappings from $S$ to $L$. We say that the pair $(\pi,\sigma)$ identifies $S$ if for each $\in S, ss$ is the least common ancestor of $\pi(s)$ and $\sigma(s)$. For $s\in S$ and $x\in V$, we say that $x$ is $s$-requested in $(\pi,\sigma)$ if $x$ lies on the path from $\pi(s)$ to $\sigma(s)$ (namely, on either of the paths from $\pi(s)$ to $s$ and from $\sigma(s)$ to $s$). We say that $x$ is requested in $(\pi,\sigma)$ if it is $s$-requested for some $s$. The pair $(\pi,\sigma)$ has unique request if every node $x$ of $T$ is requested at most once in $(\pi,\sigma)$, i.e., there exists at most one $s\in S$ such that $x$ is $s$-requested in $(\pi,\sigma)$. Note that if $(\pi,\sigma)$ has unique request then the paths from $\pi(s)$ to $\sigma(s)$ are pairwise disjoint for all the $s\in S$. We now state a few basic observations.
Remark 4.
@@
  3. (3)
    For each $\in S, ss$ is $s$-requested in $(\pi,\sigma)$.
  4. (4)
@@
Lemma 5.
Let $(\pi,\sigma)$ identifying some subset $S$ of $V\setminus L$ with unique request. Then for each $a\in\pi(S)$, the node $\pi^{-1}(a)$ is the least ancestor of $a$ which belongs to $S$.
Proof 3.1.
Let $a\in\pi(S)$ and let $s=\pi^{-1}(a)$. By definition, $s$ is an ancestor of $a$ which belongs to $S$. Let $y$ be the least ancestor of $a$ that is contained in $S$. As $s$ is an ancestor of $a$ belonging to $S$, $y$ must be a descendant of $s$. Hence, $y$ is $s$-requested in $(\pi,\sigma)$. Additionally, by item 3 of 4, $y$ is $y$-requested in $(\pi,\sigma)$. Since $(\pi,\sigma)$ has unique request, $y=s$. Thus, $s$ is the least ancestor of $a$ which belongs to $S$.
Notice that, by item 1 of 4, a similar result holds for each $b\in\sigma(S)$. It follows that the sets $\pi(S)$ and $\sigma(S)$ characterize $(\pi,\sigma)$.
Lemma 6.
Let $S\subseteq V\setminus L$, and $(\pi,\sigma)$ and $(\pi^{\prime},\sigma^{\prime})$ be two pairs of injections from $S$ to $L$ identifying $S$ with unique request. If $\pi(S)=\pi^{\prime}(S)$ and $\sigma(S)=\sigma^{\prime}(S)$, then $\pi=\pi^{\prime}$ and $\sigma=\sigma^{\prime}$.
Proof 3.2.
Let $s\in S$, $a=\pi(s)$, and $s^{\prime}=\pi^{\prime-1}(a)$. Both $s$ and $s^{\prime}$ are the least ancestor of $a$ which belongs to $S$, hence $s=s^{\prime}$ and $\pi^{\prime}(s)=a$. Thus $\pi=\pi^{\prime}$. By item 1 of 4, we also obtain that $\sigma=\sigma^{\prime}$.
Let $A$ and $B$ be two subsets of $L$ that are disjoint and of same cardinality. We call such a pair $(A,B)$ a bi-colouring. We say that $(A,B)$ identifies $S$ if there exists a pair $(\pi,\sigma)$ identifying $S$ with unique request such that $\pi(S)=A$ and $\sigma(S)=B$. By the previous lemma, for a fixed set $S\subseteq V\setminus L$ and a fixed bi-colouring $(A,B)$ of $L$, the pair $(\pi,\sigma)$ identifying $S$ with $\pi(S)=A$ and $\sigma(S)=B$ is unique when it exists. We will also prove that $S$ is actually uniquely determined from $(A,B)$. Before that, we state the following technical lemma.
Lemma 7.
@@
  1. (1)
    $\notin S, xx$ is not requested in $(\pi,\sigma)$, and for each child $c$ of $x$, $\left|A\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$;
  2. (2)
    $\notin S, xx$ is requested in $(\pi,\sigma)$, and there exists one leaf $z\in(A\cup B)\cap V(T_{x})$ such that, for each child $c$ of $x$, $\left|(A\setminus\{z\})\cap V(T_{c})\right|=\left|(B\setminus\{z\})\cap V(T_{c})\right|$;
  3. (3)
    $\in S, xx$ is requested in $(\pi,\sigma)$, and there exists two leaves $a\in A\cap V(T_{x})$ and $b\in B\cap V(T_{x})$ such that, for each child $c$ of $x$, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|(B\setminus\{b\}\cap V(T_{c}))\right|$ and $\{a,b\}\nsubseteq V(T_{c})$.
In particular, $x\in S$ if and only if there exists $a\in A\cap V(T_{x})$ and $b\in B\cap V(T_{x})$ such that, for each child $c$ of $x$, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|(B\setminus\{b\}\cap V(T_{c}))\right|$ and $\{a,b\}\nsubseteq V(T_{c})$.
For an illustration of the different cases see Figure 4 where $x_{1}$ is an example of a node satisfying the first case, $x_{2}$ an example for the second case and $x_{3}$ for the third.
Proof 3.3.
Let $x$ be an inner node. We consider the set $S^{\prime}=S\setminus(V(T_{x})\setminus\{x\})$ of all nodes which are not proper descendants of $x$ and the restrictions $\pi^{\prime}$ and $\sigma^{\prime}$ of, respectively, $\pi$ and $\sigma$ to $S^{\prime}$. By item 2 of 4 $(\pi^{\prime},\sigma^{\prime})$ identify $S^{\prime}$ with unique request. We denote $A^{\prime}=\pi^{\prime}(S^{\prime})$ and $B^{\prime}=\sigma^{\prime}(S^{\prime})$, thus $(A^{\prime},B^{\prime})$ identify $S^{\prime}$. Let $A^{\prime}_{x}=A^{\prime}\cap V(T_{x})$ and $B^{\prime}_{x}=B^{\prime}\cap V(T_{x})$ be the sets of images of nodes in $S^{\prime}$ under $\pi^{\prime}$, $\sigma^{\prime}$, respectively, contained in $T_{x}$. In case $a$ is an element of $A^{\prime}_{x}$, the element $s_{a}=\pi^{-1}(a)$ is an ancestor of $x$ because it is an ancestor of $a\in V(T_{x})$ and belongs to $S^{\prime}$ whence not to $V(T_{x})\setminus\{x\}$. Therefore, $x$ is $s_{a}$-requested. As $(\pi^{\prime},\sigma^{\prime})$ has unique request, there exists at most one element in $A^{\prime}_{x}$. Similarly, $B^{\prime}_{x}$ has size at most $1$. Moreover, $A^{\prime}_{x}\cap B^{\prime}_{x}=\emptyset$ by item 4 of 4. We thus we have three cases:
Case $A^{\prime}_{x}\cup B^{\prime}_{x}=\emptyset$:
@@
Case $A^{\prime}_{x}\cup B^{\prime}_{x}=\{a\}$:
@@
Let $c$ be a child of $x$. If $\left|A\cap V(T_{c})\right|\neq\left|B\cap V(T_{c})\right|$, then it means that there exists $z\in(A\cup B)\cap V(T_{c})$, such that either $z\in A$ and $s_{z}=\pi^{-1}(z)\notin V(T_{c})$, or $z\in B$ and $s_{z}=\sigma^{-1}(z)\notin V(T_{c})$. In both cases, $s_{z}$ is a proper ancestor of $c$, whence an ancestor of $x$. Thus, $x$ is $s_{z}$-requested in $(\pi,\sigma)$. However, $x$ is $s_{a}$-requested in $(\pi,\sigma)$ which has unique request, hence $s_{z}=s_{a}$. If $z\in B$, it follows that $z=b$, which contradicts the fact $b\notin V(T_{x})$. Hence, $z\in A$ and thus, $z=\pi(s_{a})=a$. On the other hand, if $\left|A\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$ then $a\notin V(T_{c})$. If $a\in V(T_{c})$ then $c$ would be $s_{a}$ requested and additionally (as $\left|(A\setminus\{a\}\cap V(T_{c})\right|<\left|B\cap V(T_{c})\right|$) there has to be $z\in T_{c}$, $z\neq a$ for which $s_{z}=\sigma(z)^{-1}$ is an ancestor of $c$. In particular, $s_{a}\neq s_{z}$ as $c$ is the least common ancestor of $a$ and $z$, but $s_{a}$ is a proper ancestor of $c$. Hence, $c$ is requested by both $s_{a}$ and $s_{z}$ contradicting the assumption that $(\pi,\sigma)$ has unique requests. Therefore, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$ for each child $c$ of $x$.
Case $A^{\prime}_{x}=\{a\}$ and $B^{\prime}_{x}=\{b\}$:
Let $s_{a}=\pi^{-1}(a)$ and $s_{b}=\sigma^{-1}(b)$. The node $x$ is $s_{a}$-requested and $s_{b}$-requested in $(\pi,\sigma)$, so, by the unique request property, $s_{a}=s_{b}$. Since $s_{a}$ is the least common ancestor of $a$ and $b$, it belongs to $V(T_{x})$ whence $s_{a}=x$ implying $x\in S$.
Let $c$ be a child of $x$. If $\left|A\cap V(T_{c})\right|\neq\left|B\cap V(T_{c})\right|$, then there exists $z\in(A\cup B)\cap V(T_{c})$ such that either $z\in A$ and $s_{z}=\pi^{-1}(z)\notin V(T_{c})$, or $z\in B$ and $s_{z}=\sigma^{-1}(z)\notin V(T_{c})$. In both cases, $s_{z}$ is a proper ancestor of $c$, whence an ancestor of $x$, and thus $x$ is $s_{z}$-requested in $(\pi,\sigma)$. However, $x$ is $x$-requested in $(\pi,\sigma)$ which has unique request, hence $s_{z}=x$ and thus $z\in\{a,b\}$. On the other hand, similarly to the previous case, if $\left|A\cap V(T_{c})\right|=\left|B\cap V(T_{c})\right|$, then neither $a$ nor $b$ can be contained in $V(T_{c})$. Therefore, $\left|(A\setminus\{a\})\cap V(T_{c})\right|=\left|(B\setminus\{b\})\cap V(T_{c})\right|$ for each child $c$ of $x$. Because the least common ancestor of $a$ and $b$ is $x$, there is no child $c$ of $x$ containing both $a$ and $b$ as leaves, i.e., $\{a,b\}\nsubseteq V(T_{c})$.